OpenStudy (anonymous):

How do I solve this? tan(x- 10) = cot(2x + 17)

OpenStudy (ranga):

Rewrite cot(A) as tan(pi/2 - A).

OpenStudy (anonymous):

How does that work?

OpenStudy (ranga):

tan(x- 10) = cot(2x + 17) cot(2x+17) = tan(pi/2 - 2x - 17) tan(x- 10) = tan(pi/2 - 2x - 17) x - 10 = pi/2 - 2x - 17 solve for x. And since tan has a period of pi, the general solution will be whatever you get for x above plus/minus pi.

OpenStudy (ranga):

plus/minus N*pi.

OpenStudy (anonymous):

So since they're reciprocals we can just subtract by 180

OpenStudy (ranga):

All "co" trig functions are pi/2 or 90 degree phase sifted from the corresponding trig function. sin(A) = "co"sine(pi/2 - A) "co"sine(A) = sin(pi/2 - A) sec(A) ="co"sec(pi/2-A) "co"sec(A) = sec(pi/2-A) tan(A) = "co"tan(pi/2-A) "co"tan(A) = tan(pi/2-A)

OpenStudy (ranga):

OR, sin(A) = cos(pi/2-A) cos(A) = sin(pi/2-A) sec(A) = cosec(pi/2-A) cosec(A) = sec(pi/2-A) tan(A) = cot(pi/2-A) cot(A) = tan(pi/2-A)

OpenStudy (ranga):

*phase shifted*

OpenStudy (ranga):

Also, you will get a second set of solutions by another identity: cot(A) = -tan(pi/2 + A) = tan(-pi/2 - A)

OpenStudy (anonymous):

Thank you sooo much

OpenStudy (ranga):

You are welcome.